3|x/3+1|=8 I need the steps to get to the right answer....
Found 2 solutions by Alan3354, lwsshak3:
Answer by Alan3354(69443)

3|x/3+1|=8
Divide by 3
|x/3+1| = 8/3
--------------
Case 1:
x/3+1 = 8/3
Multiply by 3
x + 3 = 8
x = 5
-------------
Case 2:
-(x/3+1) = 8/3
-x - 3 = 8
-x = 11
x = -11

Answer by lwsshak3(11628)

3|x/3+1|=8
**
3|x/3+1|=8
two conditions will satisfy this expression:
3((x/3)+1)=8
x+3=8
x=5
..
3(-(x/3)+1)=8
-x-3=8
-x=11=-11
..
ans:
x=5 or -11